Please help me solve by any method:
x=4y-3
-3x-y=0

Answer by mananth(16946)

x=4y-3
-3x-y=0
..
1..x-4y=-3
2. -3x-y=0
multiply eq 1 by 3 and add the second equation
3x-12y-3x-4y=-9
-16y=-9
y=-9/-16
y=9/16
plug the value of y in eq1
x=4*9/16 -3
x=36-48 /16
x=-12/16
x=-3/4